OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (akashdeepdeb):

If two lines are parallel, they are said to have the same slope. So, find the slope (m) of the line you are given by the equation: \(y = mx + c\) And then, since you already given a point (4, -2) on the line, use the equation of a line: \((y-y_1) = m(x-x_1)\)